# Dimensionality reduction techniques to support

### v1

@@ 2.2 Dimensionality reduction methods
$X\rightarrow Z=f_{1}(X)\rightarrow\hat{X}=f_{2}(Z)\text{ where }X,\hat{X}\in\mathbb{R}^{N,T},\ Z\in\mathbb{R}^{N,K}.$
The transformations $f_{1}$ and $f_{2}$ are such that the reconstruction error is minimized i.e.
@@
$\hat{x}_{i}(t)=g_{2}\Big{(}\sum_{k=1}^{K}g_{1}(x_{i}\cdot W_{1}(:,k))W_{2}(k,t)\Big{)}\ i=1,\ldots,N,\ t=1,\ldots,T,$ (6)
where $W_{1}\in\mathbb{R}^{T,K}$, $W_{2}\in\mathbb{R}^{K,T}$ are layers’ weight matrices, also called encoder and decoder, and $g_{1}$, $g_{2}$ are activation functions. The layers’ weight matrices are determined by common gradient descent algorithms like Adam [15].
@@
$\arg\min_{\hat{X}^{\prime}:\ rank(\hat{X}^{\prime})\leq K}||X-\hat{X}^{\prime}||_{F}^{2}$
is given by the truncated SVD. The latter approximates excellently data with linear relationships. On the other hand, concerning the reconstruction of nonlinear data, autoencoders outperform PCA, as empirical results in different fields show e.g. applications on image reconstruction. Theoretically, the difference between the problems tackled by PCA and AE is the constraint on the rank of $\hat{X}$. For PCA we impose $rank(\hat{X})\leq K$, so there are no more than $K$ columns of $\hat{X}$ which are linearly independent. Equivalently, we have no more than $K$ independent features. On the other hand, AEs aim at minimizing the reconstruction error without any constraint on $rank(\hat{X})$. This means we could end up with $rank(\hat{X})\in(K,T]$ so, more independent features.
Pros and cons of the different methods
@@
3 Data
3.1 Transaction reporting database
@@
3.2 Price sensitive events database
@@
Our data report for each PSE the stock, its date, and the time window for insider trading investigation. This period varies depending on the type of PSE, which leads to different definitions of the time at which an information starts to be considered price sensitive. In Table 1, the PSEs database is displayed.
@@
Table 1: Price sensitive events. The table reports the stock name, the date of the PSE, and the investigation period.
4 Results
As a first case study, we focus on the asset Industria Macchine Automatiche (IMA) whose takeover bid was announced on July 28, 2020. Figure 1 shows the price dynamics of this asset. The impact that the PSE had on the share price is evident: there is an increase of $13.16\%$ on the day of the announcement and the takeover bid’s price $68.0$ Euro is reached. In analogy to [19], we identify the reference period as the time window going from January 2, 2020 to June 28, 2020. Instead, the business month preceding the PSE i.e. July 28, 2020, is the investigation period, as outlined in Table 1.
@@ 4 Results
Anomaly detection with PCA
@@
After a feature scaling as pre-processing step, we run our method as illustrated in Section 2. As an example of the reconstructed trading position, in Figure 2 we show the position of an investor compared to its reconstructed counterpart obtained by PCA. Most of the days the reconstruction is quite close to the original trajectory and the anomaly detection method identifies the days and investors for which the discrepancy, i.e. the reconstruction error, is large.
@@
Applying our anomaly detection method, we obtain $1,246$ potential insiders out of the $13,225$ total investors. Given the high percentage of anomalous investors that we obtain, their ranking, following the procedure explained at the end of Section 2.1, is fundamental to provide more insight.
@@
Figure 4: PCA on IMA. Top panels. Two anomalous investors identified by PCA but not by k-means. Bottom left. A refers to the trading position computed setting $t_{0}$ = January 2, 2019. B refers to the trading position computed setting $t_{0}$ = January 2, 2020.. Bottom right. Investor detected by the method based on k-means and not by the method based on PCA. In all panels the vertical dashed line is the day corresponding to the beginning of the investigation period i.e. June 29, 2020.
@@
Another positive aspect of the new method based on a dimensionality reduction approach is its ability to detect given investors as anomalous, independently of the starting point chosen for the computation of her position (see Subsection 2.2). We observe that the condition in Equation 5 - which states when the reconstruction error of a given profile is the same of its shifted counterpart - holds for about $63\%$ of investors in our dataset, if January 2, 2020 and January 2, 2019 are two choices of zeros for the computation of the positions. In the bottom left panel of Figure 4 we show the position of one investor when setting to zero the position on January 2, 2020 (blue) or on January 2, 2019 (black). Interestingly, this investor is identified as anomalous by the method based on PCA, but not by the method based on k-means. The fact that with the choice $t_{0}=\text{January}\ 2,\ 2019$ this investor is identified as anomalous could surprise given we are focusing on detecting insider trading related to the announcement of a takeover bid. We recall that after this kind of PSE, the price increases and so, insiders are likely to have positive positions before the PSE. Indeed, this investor is not identified as anomalous by the method based on k-means given her negative value of signed turnover. On the other hand, if the profile is computed with the choice $t_{0}=\text{January}\ 2,\ 2020$, she has a net buying position before the PSE and signed turnover equal to $1$. Therefore, only with this choice of $t_{0}$, she is identified as soft discontinuous by the method based on k-means.
@@
It is also interesting to investigate why some investors are detected by the method based on k-means and not by the method based on PCA. The total number of these investors is $79$; among them, $27$ are investors with constant position in terms of shares and so, they are not included in the analysis based on PCA. The majority of the remaining profiles are of the type displayed in the bottom right part of Figure 4. They are investors who have a null position in the investigation period, if positions are computed in shares. Instead, if positions are computed in Euros, as in the method based on k-means, the signed turnover and the maximum exposure of these investors are equal to $1$ in the investigation period. They are in the best rewarding position and this makes them extremely suspicious according to the clustering approach of [19]. On the contrary, these investors are not identified by the method based on PCA, since the condition (4) relative to Equation 3 does not hold. That condition requires that the difference between the position on the PSE and the position on the first day of the reference period is larger than 0.5. For investors like PF_978820 (bottom right part of Figure 4), this difference is null since the investor closes her position before the PSE.
@@
Going nonlinear: the Autoencoder
@@
Name Neurons of the hidden layers Encoding Neurons of the hidden layers
in the encoder dimension in the decoder
@@
Bearing this in mind, we investigate the use of nonlinear autoencoders for our problem. The symmetric architectures we employ are schematized in Table 2. The number of neurons is chosen according to the geometric pyramid rule [18] and for all architectures the activation function of the hidden layers is the ReLU, while the activation function of the output layer is the hyperbolic tangent. This last choice allows to exploit the nonlinearity of the neural networks and yet to produce outputs with values in the interval $[-1,1]$, the domain of the normalized trading positions. The loss function is the mean squared error (MSE) and Adam [15] is used as optimization algorithm.
Model $||X-\hat{X}||_{F}$ EVS $\bar{s^{*}}$ $\bar{s^{*}}_{anomalous}$ $\bar{s^{*}}_{normal}$ $M_{1}$ $M_{2}$
@@
Table 3: Metrics for different dimensionality reductions of the IMA dataset. EVS is the explained variance score and it is defined as $100\Big{(}1-\frac{Var(X-\hat{X})}{Var(X)}\Big{)}$; $\bar{s^{*}}$ is the mean anomaly score; $\bar{s^{*}}_{anomalous}$ is the mean anomaly score for investors who are identified as hard/soft by the method based on k-means; $\bar{s^{*}}_{normal}$ is the mean anomaly score for all investors except the ones identified as hard/soft by the method based on k-means; $M_{1}=\frac{\bar{s^{*}}_{anomalous}-\bar{s^{*}}_{normal}}{\bar{s^{*}}_{normal}}$; $M_{2}=\frac{\bar{s^{*}}_{anomalous}}{\bar{s^{*}}}$. The results related to the autoencoders are averaged over $10$ runs.
The PCA and the $4$ autoencoder architectures of Table 2 are run on our data set of trading positions for IMA. A comparison between the reconstructed profile of an investor, obtained with PCA and a type of autoencoder is represented in Figure 5. Greater smoothness is associated to the profile obtained by AE-1 however, the overall similarity between the two different profile reconstructions is evident. Table 3 summarizes the main results for all autoencoders’ architectures in terms of several metrics. Due to their nonlinear character, we expect deep autoencoders can capture more complex features in data. This could lead to the identification of characteristics for the data compression which are more relevant than the ones identified by PCA, which is a linear model. However, we need at least $5$ hidden layers to outperform PCA in terms of the loss function $||X-\hat{X}||_{F}$. The lower MSE is accompanied by greater explained variance score (EVS) which is defined as [^15] Notice that the EVS is different from a common metric that is usually employed for PCA, that is the explained variance ratio (EVR), defined as $EVR=\frac{\sum_{k=1}^{K}var(z_{k})}{\sum_{t=1}^{T}var(x_{t})},$ where $x_{t}\in\mathbb{R}^{N}$ are the columns of $X\in\mathbb{R}^{N,T}$, that is the data matrix, and $z_{k}\in\mathbb{R}^{N}$ are the columns of $Z_{K}\in\mathbb{R}^{N,K}$, that is $X$’s representation in the $K$-dimensional latent space. The EVR is not a good indicator for autoencoders since it strictly depends on the activation functions we choose. Therefore, EVS is preferred to EVR, thus allowing a fair comparison with the PCA results.
$EVS=100\Bigg{(}1-\frac{Var(X-\hat{X})}{Var(X)}\Bigg{)}.$
@@
Nonetheless, we are not solely interested in better reconstruction errors. A larger gap between the errors of the anomalous investors and the ones who are not anomalous is desired. Given our unavailability of labels, we compare the mean anomaly score of investors who are detected as hard/soft by the method based on k-means [19] i.e. $\bar{s^{*}}_{anomalous}$, and the mean anomaly score of the other investors i.e. $\bar{s^{*}}_{normal}$.
The two metrics $M_{1}$ and $M_{2}$, that we introduce, shed light on this issue. They are defined as
$M_{1}=\frac{\bar{s^{*}}_{anomalous}-\bar{s^{*}}_{normal}}{\bar{s^{*}}_{normal}}$
and
$M_{2}=\frac{\bar{s^{*}}_{anomalous}}{\bar{s^{*}}}.$
We obtain that the model AE-1 has the greatest values of $M_{1}$ and $M_{2}$ i.e. it leads to a greater gap between the anomaly scores of our proxy of anomalous investors and the others. On the other hand, AE-4, which allows to obtain the lowest error in the data reconstruction, leads to the lowest value of the two quantities. It is important to stress that the comparison between the values of $M_{1}$ and $M_{2}$ that are obtained with different models, provide information which could be useful to our insiders detection task. However, the anomaly score is not the only quantity which determines whether an investor is identified as anomalous; the distribution of the times corresponding to the largest reconstruction errors also plays a fundamental role, both on the identification and on the ranking. We will investigate this issue in the following by employing both AE-1 and AE-4 for the anomaly detection step.
Anomaly detection with autoencoders
@@
Table 4: IMA: anomaly detection. $A$ is the set of investors identified as anomalous. $I$ is defined as $I=|A\cap A^{KM}|$ where $A_{KM}$ is set of investors identified as hard/soft by the method based on k-means. $I_{n}$ is defined as $I_{n}=|A_{n}\cap A^{KM}|$ where $A_{n}$ is the set of the first $n$ ranked anomalous investors.
We perform our anomaly detection task by employing two different architectures of autoencoders i.e AE-1 and AE-4. The main results are summarized in Table 4 and compared with the results of PCA.
@@
When we focus on the first $500$ ranked investors, PCA is the method with the largest overlap with the method based on k-means. This cannot be explained by the anomaly score values since, as shown in Table 3, the value of the metrics $M_{1}$, $M_{2}$ obtained by PCA are lower than the one obtained by AE-1. The two different dimensionality reduction methods lead to different histograms of the times corresponding to the largest reconstruction errors, which are shown in the right panels of Figures 3 and 14. While the histogram obtained with PCA has a maximum on April 17, 2020, the one obtained with AE-1 has the highest peak on the day of the PSE. Contrary to PCA, the autoencoder AE-1 is able to provide a dimensionality reduction where the trading activity on April 17, 2020 is treated as “normal” for a large fraction of investors. Therefore, the decrease of the overlapping could be ascribed to the change in the distribution of the times corresponding to the largest anomaly scores.
@@
Instead, among the first $500$ ranked investors identified by AE-1, $62$ are not identified by PCA. Among them, the investors ranked $15$ (Figure 15), $114,146,396$ are identified as hard by the method based on k-means. The others are of the type represented in the top left panel of Figure 4. In a similar way, among the first $500$ ranked identified by AE-4, investors who are not detected by PCA are $16$. Among them, $3$ are hard and are like the investor in Figure 17, the others are like the one in the top left panel of Figure 4. It is evident the ability of the autoencoders to capture as anomalous a type of profile like the ones in Figure 15 - 17, that were not identified by PCA and that are hard according to the method based on k-means. Moreover, again, the method based on dimensionality reduction approaches shows its independence of the choice of the investigation period.
@@
4.1 Other case studies

### v2

@@ II-B2 Autoencoders
$X\rightarrow Z=f_{1}(X)\rightarrow\hat{X}=f_{2}(Z)$
where $X,\hat{X}\in\mathbb{R}^{N,T},\ Z\in\mathbb{R}^{N,K}$. The transformations $f_{1}$ and $f_{2}$ are such that the reconstruction error is minimized i.e.
@@
$\hat{x}_{i}(t)=g_{2}\Big{(}\sum_{k=1}^{K}g_{1}(x_{i}\cdot W_{1}(:,k))W_{2}(k,t)\Big{)}$ (6)
for $i=1,\ldots,N,\ t=1,\ldots,T$, where $W_{1}\in\mathbb{R}^{T,K}$, $W_{2}\in\mathbb{R}^{K,T}$ are layers’ weight matrices, also called encoder and decoder, and $g_{1}$, $g_{2}$ are activation functions. The layers’ weight matrices are determined by common gradient descent algorithms like Adam [34].
@@
$\arg\min_{\hat{X}^{\prime}:\ \text{rank}(\hat{X}^{\prime})\leq K}||X-\hat{X}^{\prime}||_{F}^{2}$
is given by the truncated SVD. The latter approximates excellently data with linear relationships. On the other hand, concerning the reconstruction of nonlinear data, autoencoders outperform PCA, as empirical results in different fields show e.g. applications on image reconstruction. Theoretically, the difference between the problems tackled by PCA and AE is the constraint on the rank of $\hat{X}$. For PCA we impose $\text{rank}(\hat{X})\leq K$, so there are no more than $K$ columns of $\hat{X}$ which are linearly independent. Equivalently, we have no more than $K$ independent features. On the other hand, AEs aim at minimizing the reconstruction error without any constraint on $\text{rank}(\hat{X})$. This means we could end up with $\text{rank}(\hat{X})\in(K,T]$ so, more independent features.
II-B3 Pros and cons of the different methods
@@ II-B4 Choice of K
III Data
III-A Transaction reporting database
@@
III-B Price sensitive events database
@@
Our data report for each PSE the stock, its date, and the time window for insider trading investigation. This period varies depending on the type of PSE, which leads to different definitions of the time at which an information starts to be considered price sensitive. In Table I, the PSEs database is displayed.
TABLE I: Price sensitive events data set.
@@
IV Results
As a first case study, we focus on the asset Industria Macchine Automatiche (IMA) whose takeover bid was announced on July 28, 2020. Figure 1 shows the price dynamics of this asset. The impact that the PSE had on the share price is evident: there is an increase of $13.16\%$ on the day of the announcement and the takeover bid’s price $68.0$ Euro is reached. In analogy to [4], we identify the reference period as the time window going from January 2, 2020 to June 28, 2020. Instead, the business month preceding the PSE i.e. July 28, 2020, is the investigation period, as outlined in Table I.
@@ IV Results
IV-A Anomaly detection with PCA
@@
After a feature scaling as pre-processing step, we run our method as illustrated in Section II. As an example of the reconstructed trading position, in Figure 2 we show the position of an investor compared to its reconstructed counterpart obtained by PCA. Most of the days the reconstruction is quite close to the original trajectory and the anomaly detection method identifies the days and investors for which the discrepancy, i.e. the reconstruction error, is large.
@@
Applying our anomaly detection method, we obtain $1,246$ potential insiders out of the $13,225$ total investors. Given the high percentage of anomalous investors that we obtain, their ranking, following the procedure explained at the end of Section II-A, is fundamental to provide more insight.
@@
Figure 4: PCA on IMA. Top panels. Two anomalous investors identified by PCA but not by k-means. Bottom left. A refers to the trading position of an investor computed setting $t_{0}$ = January 2, 2019. B refers to the trading position of the same investor computed setting $t_{0}$ = January 2, 2020. Bottom right. Investor detected by the method based on k-means and not by the method based on PCA. In all panels the vertical dashed line is the day corresponding to the beginning of the investigation period i.e. June 29, 2020.
@@
Another positive aspect of the new method based on a dimensionality reduction approach is its ability to detect given investors as anomalous, independently of the starting point chosen for the computation of her position (see Subsection II-B). We observe that the condition in Equation 5 - which states when the reconstruction error of a given profile is the same of its shifted counterpart - holds for about $63\%$ of investors in our dataset, if January 2, 2020 and January 2, 2019 are two choices of zeros for the computation of the positions. In the bottom left panel of Figure 4 we show the position of one investor when setting to zero the position on January 2, 2020 (blue) or on January 2, 2019 (black). Interestingly, this investor is identified as anomalous by the method based on PCA, but not by the method based on k-means. The fact that with the choice $t_{0}=\text{January}\ 2,\ 2019$ this investor is identified as anomalous could surprise given we are focusing on detecting insider trading related to the announcement of a takeover bid. We recall that after this kind of PSE, the price increases and so, insiders are likely to have positive positions before the PSE. Indeed, this investor is not identified as anomalous by the method based on k-means given her negative value of signed turnover. On the other hand, if the profile is computed with the choice $t_{0}=\text{January}\ 2,\ 2020$, she has a net buying position before the PSE and signed turnover equal to $1$. Therefore, only with this choice of $t_{0}$, she is identified as soft discontinuous by the method based on k-means.
@@
It is also interesting to investigate why some investors are detected by the method based on k-means and not by the method based on PCA. The total number of these investors is $79$; among them, $27$ are investors with constant position in terms of shares and so, they are not included in the analysis based on PCA. The majority of the remaining profiles are of the type displayed in the bottom right part of Figure 4. They are investors who have a null position in the investigation period, if positions are computed in shares. Instead, if positions are computed in Euros, as in the method based on k-means, the signed turnover and the maximum exposure of these investors are equal to $1$ in the investigation period. They are in the best rewarding position and this makes them extremely suspicious according to the clustering approach of [4]. On the contrary, these investors are not identified by the method based on PCA, since the condition (d) relative to Equation 3 does not hold. That condition requires that the difference between the position on the PSE and the position on the first day of the reference period is larger than 0.5. For investors like PF_978820 (bottom right part of Figure 4), this difference is null since the investor closes her position before the PSE.
@@
IV-B Going nonlinear: the Autoencoder
@@
TABLE II: Autoencoders’ architectures.
Name Neurons of the Encoding Neurons of the
hidden layers dimension hidden layers
in the encoder in the decoder
@@
Bearing this in mind, we investigate the use of nonlinear autoencoders for our problem. The symmetric architectures we employ are schematized in Table II. The number of neurons is chosen according to the geometric pyramid rule [37] and for all architectures the activation function of the hidden layers is the ReLU, while the activation function of the output layer is the hyperbolic tangent. This last choice allows to exploit the nonlinearity of the neural networks and yet to produce outputs with values in the interval $[-1,1]$, the domain of the normalized trading positions. The loss function is the mean squared error (MSE) and Adam [34] is used as optimization algorithm.
TABLE III: Metrics for different dimensionality reductions of the IMA dataset. The results related to the autoencoders are averaged over $10$ runs.
Model $||X-\hat{X}||_{F}$ EVS $\bar{s^{*}}$ $\bar{s^{*}}_{anomalous}$ $\bar{s^{*}}_{normal}$ $\text{M}_{1}$ $\text{M}_{2}$
@@
The PCA and the $4$ autoencoder architectures of Table II are run on our data set of trading positions for IMA. A comparison between the reconstructed profile of an investor, obtained with PCA and a type of autoencoder is represented in Figure 5. Greater smoothness is associated to the profile obtained by AE-1 however, the overall similarity between the two different profile reconstructions is evident. Table III summarizes the main results for all autoencoders’ architectures in terms of several metrics. Due to their nonlinear character, we expect deep autoencoders can capture more complex features in data. This could lead to the identification of characteristics for the data compression which are more relevant than the ones identified by PCA, which is a linear model. However, we need at least $5$ hidden layers to outperform PCA in terms of the loss function $||X-\hat{X}||_{F}$. The lower MSE is accompanied by greater explained variance score (EVS) which is defined as [^8] Notice that the EVS is different from a common metric that is usually employed for PCA, that is the explained variance ratio (EVR), defined as $\text{EVR}=\frac{\sum_{k=1}^{K}\text{var}(z_{k})}{\sum_{t=1}^{T}\text{var}(x_{t})},$ where $x_{t}\in\mathbb{R}^{N}$ are the columns of $X\in\mathbb{R}^{N,T}$, that is the data matrix, and $z_{k}\in\mathbb{R}^{N}$ are the columns of $Z_{K}\in\mathbb{R}^{N,K}$, that is $X$’s representation in the $K$-dimensional latent space. The EVR is not a good indicator for autoencoders since it strictly depends on the activation functions we choose. Therefore, EVS is preferred to EVR, thus allowing a fair comparison with the PCA results.
$\text{EVS}=100\Bigg{(}1-\frac{\text{Var}(X-\hat{X})}{\text{Var}(X)}\Bigg{)}.$
@@
Nonetheless, we are not solely interested in better reconstruction errors. A larger gap between the errors of the anomalous investors and the ones who are not anomalous is desired. Given our unavailability of labels, we compare the mean anomaly score of investors who are detected as hard/soft by the method based on k-means i.e. $\bar{s^{*}}_{anomalous}$, and the mean anomaly score of the other investors i.e. $\bar{s^{*}}_{normal}$.
The two metrics $\text{M}_{1}$ and $\text{M}_{2}$, that we introduce, shed light on this issue. They are defined as
$\text{M}_{1}=\frac{\bar{s^{*}}_{anomalous}-\bar{s^{*}}_{normal}}{\bar{s^{*}}_{normal}},\ \text{M}_{2}=\frac{\bar{s^{*}}_{anomalous}}{\bar{s^{*}}}.$
We obtain that the model AE-1 has the greatest values of $\text{M}_{1}$ and $\text{M}_{2}$ i.e. it leads to a greater gap between the anomaly scores of our proxy of anomalous investors and the others. On the other hand, AE-4, which allows to obtain the lowest error in the data reconstruction, leads to the lowest value of the two quantities. It is important to stress that the comparison between the values of $\text{M}_{1}$ and $\text{M}_{2}$ that are obtained with different models, provide information which could be useful to our insiders detection task. However, the anomaly score is not the only quantity which determines whether an investor is identified as anomalous; the distribution of the times corresponding to the largest reconstruction errors also plays a fundamental role, both on the identification and on the ranking. We will investigate this issue in the following by employing both AE-1 and AE-4 for the anomaly detection step.
IV-C Anomaly detection with autoencoders
TABLE IV: IMA: anomaly detection. $A$ is the set of investors identified as anomalous. $I$ is defined as $I=|A\cap A^{KM}|$ where $A_{KM}$ is set of investors identified as hard/soft by the method based on k-means. $I_{n}$ is defined as $I_{n}=|A_{n}\cap A^{KM}|$ where $A_{n}$ is the set of the first $n$ ranked anomalous investors.
@@ IV-C Anomaly detection with autoencoders
We perform our anomaly detection task by employing two different architectures of autoencoders i.e AE-1 and AE-4. The main results are summarized in Table IV and compared with the results of PCA.
@@
When we focus on the first $500$ ranked investors, PCA is the method with the largest overlap with the method based on k-means. This cannot be explained by the anomaly score values since, as shown in Table III, the value of the metrics $\text{M}_{1}$, $\text{M}_{2}$ obtained by PCA are lower than the one obtained by AE-1. The two different dimensionality reduction methods lead to different histograms of the times corresponding to the largest reconstruction errors, which are shown in the right panels of Figure 3 and Figure 14 (Appendix D). While the histogram obtained with PCA has a maximum on April 17, 2020, the one obtained with AE-1 has the highest peak on the day of the PSE. Contrary to PCA, the autoencoder AE-1 is able to provide a dimensionality reduction where the trading activity on April 17, 2020 is treated as “normal” for a large fraction of investors. Therefore, the decrease of the overlapping could be ascribed to the change in the distribution of the times corresponding to the largest anomaly scores.
@@
Instead, among the first $500$ ranked investors identified by AE-1, $62$ are not identified by PCA. Among them, the investors ranked $15$ (Figure 15 in Appendix D), $114,146,396$ are identified as hard by the method based on k-means. The others are of the type represented in the top left panel of Figure 4. In a similar way, among the first $500$ ranked identified by AE-4, investors who are not detected by PCA are $16$. Among them, $3$ are hard and are like the investor in Figure 17 (Appendix D), the others are like the one in the top left panel of Figure 4. It is evident the ability of the autoencoders to capture as anomalous a type of profile like the ones in Figure 15 and 17 (Appendix D), that were not identified by PCA and that are hard according to the method based on k-means. Moreover, again, the method based on dimensionality reduction approaches shows its independence of the choice of the investigation period.
@@
IV-D Other case studies
TABLE V: Anomaly detection on all assets, obtained by employing PCA. $N$ and $T$ are the numbers of investors and days in the data set. $K$ is the encoding dimension. $A$ is the set of investors identified as anomalous. $I$ is defined as $I=|A\cap A^{KM}|$ where $A_{KM}$ is set of investors identified as hard/soft by the method based on k-means. $I_{n}$ is defined as $I_{n}=|A_{n}\cap A^{KM}|$ where $A_{n}$ is the set of the first $n$ ranked anomalous investors.